# Infinite-dimensional next-generation reservoir computing

### v1

@@ I Introduction
The idea that we just explained can be pushed all the way to considering all past lags and polynomial orders of arbitrarily high order in the polynomial kernel regression. Even though this leads to an infinite-dimensional covariate space (hence the title of the paper), the kernel regression can be explicitly and efficiently implemented using the recurrence properties of the Volterra kernel introduced in [14]. Since the Volterra kernel regression method has as covariates the left-infinite sequence of inputs and all the monomials of all degrees constructed from these inputs, this implies that, unlike NG-RC or the polynomial kernel regression, this approach is agnostic with respect to the number of lags and the order of monomials. Moreover, the Volterra kernel has been shown in [14] to be universal in the space of fading memory input/output systems with uniformly bounded inputs, and moreover, has a computational complexity that outperforms the NG-RC whenever higher-order monomials and lags are required for modeling more functionally complex systems.
The last part of the paper contains various numerical simulations that illustrate that (i) using the more computationally tractable polynomial kernel regression, one has access to a broader feature space, which allows learning more complex systems than those typically presented when using the NG-RC methodology and (ii) the Volterra kernel is a useful tool that can produce more accurate forecasts because it allows one to take infinite lags and monomial orders into account, which is of relevance in the modeling of long-memory phenomena that exhibit functionally complex dependencies.
@@ II.3 Kernel methods
By the bounded inverse theorem, $L^{-1}$, which is linear, is bounded. By the Riez representation theorem, the adjoints of $L$ and $L^{-1}$, denoted by $L^{*}$ and $(L^{-1})^{*}$ are well defined. By (3), for $f\in\mathbb{H}_{1}$, there exists $w\in H_{1}$ such that $f(\mathbf{z})=\left<w,\Phi_{1}(\mathbf{z})\right>_{H_{1}}$ for any $\mathbf{z}\in\mathcal{Z}$. Then it is easy to see that for any $\mathbf{z}\in\mathcal{Z}$, $f(\mathbf{z})=\left<(L^{-1})^{*}w,\Phi_{2}(\mathbf{z})\right>_{H_{2}}$ so that $\mathbb{H}_{1}\subset\mathbb{H}_{2}$. The reverse inclusion is similar. ∎
@@
Note that in the framework discussed in Section II.2, where the behavior of outputs is determined by causal functionals or CCIMs, whenever the corresponding functional is continuous, and the inputs are defined on a compact input space, kernel universality implies that the elements of the corresponding RKHS can be used to uniformly approximate the data generating functional.
@@
$\bm{\alpha}^{*}=\left(\mathsf{K}+\lambda_{\textup{reg}}\mathbb{I}\right)^{-1}\mathsf{Y}_{n}.$ (8)
@@ II.4 The NG-RC methodology
$\mathbf{w}^{*}={(\mathsf{X}^{\top}\mathsf{X}+\lambda_{\textup{reg}}\mathbb{I})}^{-1}\mathsf{X}^{\top}\mathsf{Y}$ (16)
@@ III Kernelization of NG-RC
$K^{\textup{NG-RC}}(\underline{\mathbf{z}}^{\tau},\underline{\mathbf{z}^{\prime}}^{\tau})\coloneqq\Phi(\underline{\mathbf{z}}^{\tau})^{\top}\Phi(\underline{\mathbf{z}^{\prime}}^{\tau})$
@@ IV Infinite-dimensional NG-RC and Volterra kernels
$\sum_{k=0}^{p}\underbrace{\sum_{\begin{subarray}{c}k_{0}+\ldots+k_{\tau-1}=k\\ k_{0},\ldots,k_{\tau-1}\geq 0\end{subarray}}\binom{p}{p-k,k_{0},\ldots,k_{\tau-1}}\prod_{t=0}^{\tau-1}(\mathbf{z}_{i-t}^{\top}\mathbf{z}_{j-t})^{k_{t}}}_{(*)}$
@@
$K^{\textup{Volt}}(\underline{\mathbf{z}_{i}},\underline{\mathbf{z}_{j}})=1+\sum_{\tau=1}^{\infty}\lambda^{2\tau}\prod_{t=0}^{\tau-1}\sum_{k=0}^{\infty}\left(\theta^{2}\mathbf{z}_{i-t}^{\top}\mathbf{z}_{j-t}\right)^{k}\\ =1+\sum_{\tau=1}^{\infty}\sum_{k=0}^{\infty}\lambda^{2\tau}\theta^{2k}\underbrace{\sum_{\begin{subarray}{c}k_{0}+\ldots+k_{\tau-1}=k\\ k_{0},\ldots,k_{\tau-1}\geq 0\end{subarray}}\prod_{t=0}^{\tau-1}(\mathbf{z}_{i-t}^{\top}\mathbf{z}_{j-t})^{k_{t}}}_{(*)}$
@@
$\mathsf{K}_{i,j}=1+\frac{\lambda^{2}\mathsf{K}^{\textup{Volt}}_{i-1,j-1}}{1-\theta^{2}\left<\mathbf{z}_{i},\mathbf{z}_{j}\right>}$ (25)
@@ V.1 Data generating processes and experimental setup
For the Lorenz and Mackey-Glass dynamical systems, the task consisted of performing the usual path-continuation. During training, inputs are the spatial coordinates at time $t$ and outputs are the $(t+1)$-th spatial coordinate, and estimators are trained on a collection of $n$ input/outputs $\{(\mathbf{z}_{t},\mathbf{z}_{t+1})\}_{t=1}^{n}$. To test their performance, the estimators are run autonomusly. That is, after seeing initial input $\mathbf{z}_{n+1}$,the outputs $\widehat{\mathbf{z}}_{n+2},\ldots,\widehat{\mathbf{z}}_{n+h}$, for some forecasting horizon $h$, are fedback into the estimator as inputs. These outputs $\widehat{\mathbf{z}}_{n+j}$ for $j=2,\ldots,h+1$ are compared against the reserved set of testing values $\mathbf{z}_{n+j}$ for $j=2,\ldots,h+1$, unseen by the estimators.
@@
with the initial condition function being the constant function $z(t)=1.2$. To numerically solve this DDE, the delay interval was discretized with time step of 0.02, then the usual RK45 procedure was performed on the discretized version of the system. The resulting dataset was flattened back into a one-dimensional dataset and to reduce the size of dataset, the dataset was further spliced to take every 50-th datapoint. The final dataset consisted of 7650 points and the first 3000 points were reserved for training. The remaining points were reserved to compare against the path-continued outputs of each estimator. Due to the discretization process, the differential equation becomes a system of equations where each $z_{t}$ is a function of past observations, as is assumed by our premise in Section II.2.
@@
$H_{t}$ $=CC^{\top}+A\mathbf{r}_{t-1}\mathbf{r}_{t-1}^{\top}A^{\top}+BH_{t-1}B^{\top}$
where the input innovations $\mathbf{z}_{t}$ are Gaussian IID, and the output observations are the conditional covariances $H_{t}$. The diagonal BEKK specification is chosen where $C$ is an upper-triangular matrix, and $A$ and $B$ are diagonal matrices of dimension $d$. It is known that there exists a unique stationary and ergodic solution whenever $A_{ii}>0,|B_{ii}|<1$ for $i=1,\ldots,d$, which expresses $H_{t}$ as a highly nonlinear function of its past inputs [6] (as per our premise in Section II.2). Since the covariance matrices are symmetric, we only need to learn the outputs $\mathbf{h}_{t}=\textup{vech}(H_{t})\in\mathbb{R}^{q}$, $q\coloneqq d(d+1)/2$, where the vech operator stacks the columns of a given square matrix from the principal diagonal downwards. An existing dataset from [14] was used with input dimensions of 15 and output dimensions of 120. The dataset consisted of 3760 input and output points, the first 3007 were reserved for training and the remaining 753 were reserved for testing. Since the output points were very small, to minimize loss of accuracy due to computational truncation errors, the output values were scaled by 1000. The training output data was further normalized so that each dimension would have 0 mean and variance of 1.
Since NG-RC methodology does not typically require normalization, the NG-RC datasets were not normalized. For the polynomial kernel, kernel values could become too large and result in truncation inaccuracies, so the training inputs were normalized to have maximum of 1 and minimum of 0. Due to the construction of the Volterra kernel, the input sequence space into the kernel for a finite sample is truncated with zeros. We thus demean the training input data. Moreover, to avoid incurring truncation errors for $\theta$ and $\lambda$ values, the maximum Euclidean norm of the inputs $M$ is set to 1, by scaling the training input values. Note that for all estimators, normalization is always performed based only on information from the training values, then the testing data is shifted and scaled based on what was used for the training data. This prevents leakage of information such as the mean, standard deviation, maximum, minimum etc. to the testing datasets.
@@ V.3 Cross-validation
Note that the washout was not cross-validated for. For both NG-RC and polynomial kernel regression, washout is the number of delays taken. For Volterra kernel, a longer washout is needed to wash the effect of truncating input samples with zeros. A washout of 100 was sufficient to generate meaningful results both for the full training set and when the training sets were restricted during cross-validation.
For the path-continuation tasks (Lorenz and Mackey-Glass), to select hyperparameters, cross-validation was performed by splitting each training set into training-validation folds which overlapped. During validation, path-continuation was performed and compared with the validation set. That is, the outputs of each estimator were fed back as inputs, and these autonomusly generated outputs were compared with the validation set. With overlapping datasets, a smaller training set would be sufficient to create multiple training folds starting from different initial points, such that in each training fold, the estimator has sufficient time to capture dynamics during training. Then during the validation phase for each fold, for a good estimator, there would be dynamical evolution in the outputs generated by the estimator. For example, the estimator did not just fit the average. This leads to meaningful validation set errors which improves the selection process for optimal hyperparameters.
For the BEKK input/output forecasting task, the usual time-series training-validation folds were used. That is, the training dataset was split into equally sized sets where the $i$-th training fold was the concatenation of the first $i$ sets and the $i$-th validation fold was $(i+1)$-set. For input/output forecasting, where estimator sees, during forecasting, a new set of inputs, this method of cross-validating turned out to be sufficient for estimators to capture the dynamics of input and output variables. Note that cross-validation training and testing was made to mimic as closely as possible the actual task to be performed on the full training set, so normalization in each fold was also performed as would have been done on the full training dataset.
@@ V.4 Results
$R^{2}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{d}\sum_{u=1}^{d}\left(1-\frac{\sum_{i=1}^{h}(y_{u,i}-\widehat{y}_{u,i})^{2}}{\sum_{i=1}^{h}(y_{u,i}-\bar{y}_{u,i})^{2}}\right)$
@@
$\textup{PSDE}(\textup{PSD},\widehat{\textup{PSD}})=\sum_{u=1}^{d}\sum_{f=1}^{F}\frac{|\textup{PSD}_{u,f}-\widehat{\textup{PSD}}_{u,f}|}{\textup{PSD}_{u,f}}$
@@
BEKK NG-RC - 0.00125 0.150 0.0204 0.0166 0.644 0.125 1.565 0.332
Polnomial - 0.00124 0.150 0.0204 0.0166 0.642 0.123 1.140 0.337
Volterra - 0.00102 0.136 0.0170 0.0139 0.634 0.382 0.963 0.319
In more complex systems such as Mackey-Glass and BEKK, the Volterra reservoir and polynomial kernel regressions easily outperform the NG-RC because they have access to much richer feature spaces. In second-order systems such as the Lorenz system, even though pointwise, errors perform better than in the polynomial and Volterra kernel regressions, the climate metrics indicate that the NG-RC better captured the climate of the true Lorenz dynamical system. It could be that a lower order system offered by the NG-RC acts as a better proxy for lower order true dynamical systems, which accounts for the difference in climate performance. This difference in climate replication performance, however, is only slight. Observe that in Figure 1(f), all estimators capture the power spectral density of the original system well, even if the Volterra kernel is slightly outperformed by the polynomial kernel. For even more complex systems, both the kernel regression methods can capture the climate of the true dynamical system but the same cannot be said for NG-RC. A similar story holds when one considers the Wasserstein-1 distance. The distributions are in Figure 2(f). Even though the Wasserstein-1 distance for Volterra performs the poorest, the difference in distribution performance is still small, and the bulk of the distribution is still replicated. On the other hand, for complex systems such as the BEKK, the polynomial kernel and the NG-RC fail to replicate the climate of the original system completely.

### v2

@@ I Introduction
The idea that we just explained can be pushed all the way to considering all past lags and polynomial orders of arbitrarily high order in the polynomial kernel regression. Even though this leads to an infinite-dimensional covariate space (hence the title of the paper), the kernel regression can be explicitly and efficiently implemented using the recurrence properties of the Volterra kernel introduced in [14]. Since the Volterra kernel regression method has as covariates the left-infinite sequence of inputs and all the monomials of all degrees constructed from these inputs, this implies that, unlike NG-RC or the polynomial kernel regression, this approach is agnostic with respect to the number of lags and the order of monomials. Moreover, the Volterra kernel has been shown in [14] to be universal in the space of fading memory input/output systems with uniformly bounded inputs and, moreover, has a computational complexity that outperforms the NG-RC whenever higher-order monomials and lags are required for modeling more functionally complex systems.
The last part of the paper contains various numerical simulations that illustrate that (i) using the more computationally tractable polynomial kernel regression, one has access to a broader feature space, which allows learning more complex systems than those typically presented when using the NG-RC methodology and (ii) the Volterra kernel is a useful tool that can produce more accurate forecasts because it allows one to take infinite lags and monomial orders into account, which is of relevance in the modeling of long-memory phenomena that exhibit functionally complex dependencies. All the codes and data necessary to reproduce the results in the paper are available at https://github.com/Learning-of-Dynamic-Processes/kernelngrcvolterra.
@@ II.3 Kernel methods
By the bounded inverse theorem, $L^{-1}$, which is linear, is bounded. By the Riesz representation theorem, the adjoints of $L$ and $L^{-1}$, denoted by $L^{*}$ and $(L^{-1})^{*}$ are well defined. By (3), for $f\in\mathbb{H}_{1}$, there exists $w\in H_{1}$ such that $f(\mathbf{z})=\left<w,\Phi_{1}(\mathbf{z})\right>_{H_{1}}$ for any $\mathbf{z}\in\mathcal{Z}$. Then it is easy to see that for any $\mathbf{z}\in\mathcal{Z}$, $f(\mathbf{z})=\left<(L^{-1})^{*}w,\Phi_{2}(\mathbf{z})\right>_{H_{2}}$ so that $\mathbb{H}_{1}\subset\mathbb{H}_{2}$. The reverse inclusion is similar. ∎
@@
Note that in the framework discussed in Section II.2, where the behavior of outputs is determined by causal functionals or CCIMs, whenever the corresponding functional is continuous and the inputs are defined on a compact input space, kernel universality implies that the elements of the corresponding RKHS can be used to uniformly approximate the data generating functional.
@@
$\bm{\alpha}^{*}=\left(\mathsf{K}^{\top}\mathsf{K}+\lambda_{\textup{reg}}\mathsf{K}\right)^{-1}\mathsf{K}\mathsf{Y}_{n}.$ (8)
@@ II.4 The NG-RC methodology
$\mathbf{w}^{*}={(\mathsf{X}^{\top}\mathsf{X}+\lambda_{\textup{reg}}\mathbb{I})}^{-1}\mathsf{X}^{\top}\mathsf{Y},$ (16)
@@ III Kernelization of NG-RC
$K^{\textup{NG-RC}}(\underline{\mathbf{z}}^{\tau},\underline{\mathbf{z}^{\prime}}^{\tau})\coloneqq\Phi(\underline{\mathbf{z}}^{\tau})^{\top}\Phi(\underline{\mathbf{z}^{\prime}}^{\tau}),$
@@ IV Infinite-dimensional NG-RC and Volterra kernels
$\sum_{k=0}^{p}\underbrace{\sum_{\begin{subarray}{c}k_{0}+\ldots+k_{\tau-1}=k\\ k_{0},\ldots,k_{\tau-1}\geq 0\end{subarray}}\binom{p}{p-k,k_{0},\ldots,k_{\tau-1}}\prod_{t=0}^{\tau-1}(\mathbf{z}_{i-t}^{\top}\mathbf{z}_{j-t})^{k_{t}}}_{(*)},$
@@
$K^{\textup{Volt}}(\underline{\mathbf{z}_{i}},\underline{\mathbf{z}_{j}})=1+\sum_{\tau=1}^{\infty}\lambda^{2\tau}\prod_{t=0}^{\tau-1}\sum_{k=0}^{\infty}\left(\theta^{2}\mathbf{z}_{i-t}^{\top}\mathbf{z}_{j-t}\right)^{k}\\ =1+\sum_{\tau=1}^{\infty}\sum_{k=0}^{\infty}\lambda^{2\tau}\theta^{2k}\underbrace{\sum_{\begin{subarray}{c}k_{0}+\ldots+k_{\tau-1}=k\\ k_{0},\ldots,k_{\tau-1}\geq 0\end{subarray}}\prod_{t=0}^{\tau-1}(\mathbf{z}_{i-t}^{\top}\mathbf{z}_{j-t})^{k_{t}}}_{(*)},$
@@
$\mathsf{K}_{i,j}=1+\frac{\lambda^{2}\mathsf{K}^{\textup{Volt}}_{i-1,j-1}}{1-\theta^{2}\left<\mathbf{z}_{i},\mathbf{z}_{j}\right>},$ (25)
@@ V.1 Data generating processes and experimental setup
For the Lorenz and Mackey-Glass dynamical systems, the task consisted of performing the usual path-continuation. During training, inputs are the spatial coordinates at time $t$ and outputs are the $(t+1)$-th spatial coordinate, and estimators are trained on a collection of $n$ input/outputs $\{(\mathbf{z}_{t},\mathbf{z}_{t+1})\}_{t=1}^{n}$. To test their performance, the estimators are run autonomously. That is, after seeing initial input $\mathbf{z}_{n+1}$,the outputs $\widehat{\mathbf{z}}_{n+2},\ldots,\widehat{\mathbf{z}}_{n+h}$, for some forecasting horizon $h$, are fedback into the estimator as inputs. These outputs $\widehat{\mathbf{z}}_{n+j}$ for $j=2,\ldots,h+1$ are compared against the reserved set of testing values $\mathbf{z}_{n+j}$ for $j=2,\ldots,h+1$, unseen by the estimators.
@@
with the initial condition function being the constant function $z(t)=1.2$. To numerically solve this DDE, the delay interval was discretized with time step of 0.02, then the usual RK45 procedure was performed on the discretized version of the system. The resulting dataset was flattened back into a one-dimensional dataset, and to reduce the size of the dataset, the dataset was further spliced to take every 50-th data point. The final dataset consisted of 7650 points, and the first 3000 points were reserved for training. The remaining points were reserved to compare against the path-continued outputs of each estimator. Due to the discretization process, the differential equation becomes a system of equations where each $z_{t}$ is a function of past observations, as is assumed by our premise in Section II.2.
@@
$H_{t}$ $=CC^{\top}+A\mathbf{r}_{t-1}\mathbf{r}_{t-1}^{\top}A^{\top}+BH_{t-1}B^{\top},$
where the input innovations $\mathbf{z}_{t}$ are Gaussian IID, and the output observations are the conditional covariances $H_{t}$. The diagonal BEKK specification is chosen where $C$ is an upper-triangular matrix, and $A$ and $B$ are diagonal matrices of dimension $d$. It is known that there exists a unique stationary and ergodic solution whenever $A_{ii}>0,|B_{ii}|<1$ for $i=1,\ldots,d$, which expresses $H_{t}$ as a highly nonlinear function of its past inputs [6] (as per our premise in Section II.2). Since the covariance matrices are symmetric, we only need to learn the outputs $\mathbf{h}_{t}=\textup{vech}(H_{t})\in\mathbb{R}^{q}$, $q\coloneqq d(d+1)/2$, where the vech operator stacks the columns of a given square matrix from the principal diagonal downwards. An existing dataset from [14] was used with input dimensions of 15 and output dimensions of 120. The dataset consisted of 3760 input and output points, the first 3007 were reserved for training, and the remaining 753 were reserved for testing. Since the output points were very small, to minimize loss of accuracy due to computational truncation errors, the output values were scaled by 1000. The training output data was further normalized so that each dimension would have 0 mean and variance of 1.
Since NG-RC methodology does not typically require normalization, the NG-RC datasets were not normalized. For the polynomial kernel, kernel values could become too large and result in truncation inaccuracies, so the training inputs were normalized to have a maximum of 1 and a minimum of 0. Due to the construction of the Volterra kernel, the input sequence space into the kernel for a finite sample is truncated with zeros. We thus demean the training input data. Moreover, to avoid incurring truncation errors for $\theta$ and $\lambda$ values, the maximum Euclidean norm of the inputs $M$ is set to 1, by scaling the training input values. Note that for all estimators, normalization is always performed based only on information from the training values, then the testing data is shifted and scaled based on what was used for the training data. This prevents leakage of information such as the mean, standard deviation, maximum, minimum, etc., to the testing datasets.
@@ V.3 Cross-validation
Note that the washout was not cross-validated for. For both NG-RC and polynomial kernel regression, washout is the number of delays taken. For the Volterra kernel, a longer washout is needed to wash the effect of truncating input samples with zeros. A washout of 100 was sufficient to generate meaningful results both for the full training set and when the training sets were restricted during cross-validation.
For the path-continuation tasks (Lorenz and Mackey-Glass), to select hyperparameters, cross-validation was performed by splitting each training set into training-validation folds that overlapped. During validation, path-continuation was performed and compared with the validation set. That is, the outputs of each estimator were fed back as inputs, and these autonomously generated outputs were compared with the validation set. With overlapping datasets, a smaller training set would be sufficient to create multiple training folds starting from different initial points, such that in each training fold, the estimator has sufficient time to capture dynamics during training. Then during the validation phase for each fold, for a good estimator, there would be dynamical evolution in the outputs generated by the estimator. For example, the estimator did not just fit the average. This leads to meaningful validation set errors which improves the selection process for optimal hyperparameters.
For the BEKK input/output forecasting task, the usual time-series training-validation folds were used. That is, the training dataset was split into equally sized sets where the $i$-th training fold was the concatenation of the first $i$ sets and the $i$-th validation fold was $(i+1)$-set. For input/output forecasting, where estimator sees, during forecasting, a new set of inputs, this method of cross-validating turned out to be sufficient for estimators to capture the dynamics of input and output variables. Note that cross-validation training and testing were made to mimic as closely as possible the actual task to be performed on the full training set, so normalization in each fold was also performed as would have been done on the full training dataset.
@@ V.4 Results
$R^{2}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{d}\sum_{u=1}^{d}\left(1-\frac{\sum_{i=1}^{h}(y_{u,i}-\widehat{y}_{u,i})^{2}}{\sum_{i=1}^{h}(y_{u,i}-\bar{y}_{u,i})^{2}}\right),$
@@
$\textup{PSDE}(\textup{PSD},\widehat{\textup{PSD}})=\sum_{u=1}^{d}\sum_{f=1}^{F}\frac{|\textup{PSD}_{u,f}-\widehat{\textup{PSD}}_{u,f}|}{\textup{PSD}_{u,f}},$
@@
BEKK NG-RC 0.00125 0.150 0.0204 0.0166 0.644 0.125 1.565 0.332
Polynomial 0.00124 0.150 0.0204 0.0166 0.642 0.123 1.140 0.337
Volterra 0.00102 0.136 0.0170 0.0139 0.634 0.382 0.963 0.319
In more complex systems such as Mackey-Glass and BEKK, the Volterra reservoir and polynomial kernel regressions easily outperform the NG-RC because they have access to much richer feature spaces. In second-order systems such as the Lorenz system, even though pointwise errors perform better than in the polynomial and Volterra kernel regressions, the climate metrics indicate that the NG-RC better captured the climate of the true Lorenz dynamical system. It could be that a lower-order system offered by the NG-RC acts as a better proxy for lower-order true dynamical systems, which accounts for the difference in climate performance. This difference in climate replication performance, however, is only slight. Observe that in Figure 1(f), all estimators capture the power spectral density of the original system well, even if the Volterra kernel is slightly outperformed by the polynomial kernel. For even more complex systems, both the kernel regression methods can capture the climate of the true dynamical system but the same cannot be said for NG-RC. A similar story holds when one considers the Wasserstein-1 distance. The distributions are in Figure 2(f). Even though the Wasserstein-1 distance for Volterra performs the poorest, the difference in distribution performance is still small, and the bulk of the distribution is still replicated. On the other hand, for complex systems such as the BEKK, the polynomial kernel and the NG-RC fail to replicate the climate of the original system completely.
@@
Acknowledgments: The authors thank Daniel Gauthier for insightful discussions about the relation between NG-RC and the results in this paper. LG and JPO thank the hospitality of the Nanyang Technological University and the University of St. Gallen, respectively; it is during respective visits to these two institutions that some of the results in this paper were obtained. HLJT is funded by a Nanyang President’s Graduate Scholarship of Nanyang Technological University. JPO acknowledges partial financial support from the School of Physical and Mathematical Sciences of the Nanyang Technological University.